OpenStudy (anonymous):

1) write an equation of the line with the given slope and y-intercept. m= 8/3; (0,-4) 2) Write the slope-intercept form of the equation for the line passing through the given pair of points. (-9, 7) and (-7, 4)

OpenStudy (anonymous):

for 1) use: \[y-y_1=m(x-x_1)\] i.e. \[y-(-4)=\frac{8}{3}(x-0)\] Sorry, I can't remember what slope-intercept form is for the second question.

OpenStudy (anonymous):

ok, no worries. Thanks for your help